Finance

NST headline, "Covid-19: How to Improve Our Life." The title of this newspaper headline refers to an economic plan proposed by a minister. Minister has proposed giving every newborn baby a RM1,000.00 government savings account at birth, followed by five annual contributions of RM500.00 each. If the funds are left untouched in an investment account, then by the time each baby reaches age 65, his or her RM3,500.00 contribution will have grown to RM600,000.00 over the years, even at medium returns for a thrift-savings plan. At about 9.4%
compounded annually, the balance would grow to be RM1,005,132.00. (How would you calculate this number?) Since about 4 million babies are born each year, the proposal would cost the federal government RM4 billion annually. About 90% of the total annual tax collections of more than RM300 billion is used to pay current beneficiaries. The remaining 10% is invested in interest-bearing government bonds that finance the day-to-day expenses of the federal government. Discuss the economics of minister savings plan.
Answer #1

The Minister's economics seems to be a sound plan.

As per the plan, the amount to be invested from the year zero to 5 are given below with the compounded figures at the end of 5th year.

Year Amount. Compounded (@9.4%)amount after 5 years

0. 1000. 1000×(1.094)^5=1567.0635

1. 500. 500×(1.094)^4= 716.2082

2. 500. 500×(1.094)^3= 654.6693

3. 500. 500×(1.094)^2=598.418

4. 500. 500×(1.094)^1=547

5. 500. 500(1.094^0 = 500

Total. = 4583.359

Now, this amount of 4583.359 is to be compounded at 9.4% for next 60 years.

So, 4583.359×(1.094)^60=1005132

This is the calculation of the said amount by minister.

This is sounding good because the annual spending for this would be 4 billion which is feasible with the tax collection.

​​​​